two good dice are rolled simultaneously. let a denote the event “ the sum shown is 8" and b the event “the two show the same number". find p(a), p(b), p(a ∩ b), and p(a ∪ b).

1
Expert's answer
2022-06-02T12:40:51-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 2 & 3 & 4 & 5 & 6 & 7\\
 \hdashline
 2 & 3 & 4 & 5 & 6 & 7 & 8\\
 \hdashline
 3 & 4 & 5 & 6 & 7 & 8 & 9\\
 \hdashline
 4 & 5 & 6 & 7 & 8 & 9 & 10\\
 \hdashline
 5 & 6 & 7 & 8 & 9 & 10 & 11\\
 \hdashline
 6 & 7 & 8 & 9 & 10 & 11 & 12\\
 \hdashline
\end{array}"

The sum shown is 8: 2+6=3+5=4+4=5+3=6+2

The same number: "(1,1), (2,2), (3,3), (4,4), (5,5),(6,6)"


a)

"P(A)=\dfrac{5}{36}"

b)

"P(B)=\dfrac{6}{36}=\dfrac{1}{6}"

c)


"PA\cap B)=\dfrac{1}{36}"

d)


"PA\cup B)=\dfrac{10}{36}=\dfrac{5}{18}"